OpenStudy (anonymous):

A clothing store owner is experimenting with the layout of a clothing rack on the floor. One day, she kept track of the number of customers in the store who visited the rack of new clothing items. She found that sixty-two percent of her customers looked through that rack. The next day, she moved the rack of new items to a different location on the floor. Here's what she found: Total customers: 75 Number of customers looking at new items: 60 How did the movement of the rack affect the number of customers looking at the new items?

OpenStudy (anonymous):

compare \[\frac{60}{75}\] to 62%

OpenStudy (anonymous):

60/75 = 80% 80% > 62% The new location is better.

OpenStudy (anonymous):

Does It decrease by 2 percent, decrease by 8 percent, Increase by thirteen percent, or increase by eighteen percent...??

OpenStudy (anonymous):

this question is ambiguous

OpenStudy (anonymous):

the actual increase is 18% but be careful.

OpenStudy (anonymous):

in truth the actual "increase" is about 29% since \[\frac{18}{62}=.29\] but that is probably not what they want you to say, so stick with 18